OpenStudy (anonymous):

Suppose that GCD(a,b) = 1 and that a divides bc. Show that a must divide c.

OpenStudy (anonymous):

Suppose a|c and b|c. Then c = am and c = bn. Suppose that gcd(a,b) = 1. We may write it as a linear combination: 1 = ax + by Observe that 1 = ax + by c = acx + bcy c = a(bn)x + b(am)y c = ab(nx) + ab(my) c = ab(nx + my).

OpenStudy (anonymous):

hopes this helps

OpenStudy (anonymous):

Why does c = am and c = bm?

OpenStudy (anonymous):

And why did we suppose that a|c and b|c?